OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 625 and a2 = –125?

OpenStudy (anonymous):

first the common ratio using r= a2/a1 then use a7=a1(r)^7-1

OpenStudy (anonymous):

u should get1/(625)^2

OpenStudy (anonymous):

you would get 0.04